What is the speed of the ball after the collision?

A 120g rubber ball is thrown with a speed of 10 m/s at an oncoming 900kg car which is approaching at 30m/s, and undergoes a one-dimensional elastic collision with acar. What is the speed of the ball after the collision?
Answer #1
Let the velocity of ball be v and the car be V
elastic coefficient restitution,

e = (10 + 30)/(v- V) = 1

v-V = 40

conservation of linear momentum,

?P = 0

mv + MV - mv' - MV' = 0

-m * 10 + M * 30 - mv - M(V) = 0

all terms divided by M,

-m*10/M +30 -mv/M - (V) =0

M >>> m

mv'/M = 0

=>V = 30m/s

v= 30+40 = 70m/sec
answered by: Kalyn